# Strong coupling non-Markovian quantum thermodynamics of a finite-bath system

### v2

@@ I Introduction
In this work, we study the strong-coupling non-Markovian quantum thermodynamics of a central spin interacting with a spin bath [73, 74, 75, 76, 77]. The baths can be broadly placed into the following categories [78]: (a) Bosonic, (b) fermionic, and (c) spin baths. The spin-boson model [8] and the Caldeira-Leggett model [11] are well-known examples of bosonic baths. The dynamics of these types of models have been studied in the literature [1, 2, 3]. Conducting electrons in the Anderson and the Kondo models are examples of a fermionic bath system. Numerically exact methods, for example, the fermionic hierarchical equation of motion, have been used to obtain the dynamics of the above systems [79, 80]. Further, exact dynamics of noninteracting fermions (bosons) linearly coupled to thermal environments of noninteracting fermions can be found in [81]. Conversely, when dealing with spin baths [82], frequent use is made of perturbative methods or time-dependent quantum master equations [83, 84, 85]. Dissipative equation of motion approach has also been used to find exact dynamics of the systems surrounded by spin bath [86]. Studying these systems is crucial for understanding magnetic systems [87], quantum spin glasses [88], quantum batteries [89], NV-center [90], and superconducting systems [83], among others. Moreover, the finite nature of the bath renders the exact dynamics of the central spin model analytically tractable across arbitrary system-bath coupling regimes. This unique feature provides profound advantages for probing the non-Markovian effects on quantum thermodynamics, particularly in strong coupling regimes.
@@
Here, we develop the exact dynamics of the central spin model. This is then used to understand the quantum thermodynamic properties of the central spin. To this effect, we calculate the thermodynamic quantities of this system using a number of methods outlined above. We begin by analyzing entropy production, heat, and work while discussing the first and second laws of thermodynamics. We also find out the cost of turning the system and bath interactions on or off, and we explore a variation of the second law in finite bath models. Next, we calculate the thermal equilibrium state and entropy under strong coupling using the HMF technique, which aids in examining the third law of quantum thermodynamics. We then discuss the use of the canonical Hamiltonian as the replacement for the system’s bare Hamiltonian to calculate the thermodynamic quantities. We aim to arrive at the correct forms of the quantities, such as heat, work, internal energy, entropy production, and thermal equilibrium state in strongly coupled non-Markovian quantum systems and the correct forms of the corresponding laws of quantum thermodynamics, taking the central spin model as an example. Finally, we demonstrate the central spin model as a quantum battery, highlighting how a finite spin bath uniquely charges the central spin quantum battery. To summarize, we first present a brief discussion of the tools for handling strong-coupling non-Markovian quantum thermodynamics (Sec. II). This is followed in the later sections with new results on the exact dynamics of the central spin model and its quantum thermodynamics in the strong coupling non-Markovian regime, by applying the above tools.
@@ II Tools for exploring non-Markovian strong-coupling quantum thermodynamics
Here, we briefly dive into the key tools for probing quantum thermodynamics in a strongly coupled non-Markovian quantum system. We initiate the discussion from entropy production for a general quantum system surrounded by a bath (using this, we discuss the second law of quantum thermodynamics), followed by the discussion of the change in the system’s and bath’s internal energy and the work done as the mismatch between them. This leads us to the first law of quantum thermodynamics in a general system-bath scenario. In the next step, we discuss how work done is related to the cost of turning the system-bath interactions on or off. To analyze the thermal equilibrium between the system and the bath in the strong-coupling regime, we discuss the formalism of Hamiltonian of mean force and the mean force Gibbs state. Further, we discuss the canonical Hamiltonian of a system.
@@ II.1 Entropy production, heat and work
Considering the global system-bath dynamics for arbitrary initial states of the system and the bath [Eq. (2)], a formulation for the entropy production was developed in [91, 28]. Here, the dynamical evolution of the system is obtained by tracing out the bath, which is the origin of the irreversibility. The entropy production accounts for the contribution of two processes, one where we discard any information stored locally in the state of the bath and the other where we discard the non-local information shared between the system and the bath, and is given by
@@
[…]
@@ II.3 The theory of canonical Hamiltonian
on the Lindblad (jump) and Hamiltonian operators, respectively. It leaves the structure $-i[H_{S}(t),\rho_{S}]+\mathcal{D}_{t}(\rho_{S})$ intact. To get rid of this ambiguity and to render a unique dissipator, the jump operators $L_{j}^{\prime}$s need to be trace-less, as shown in [98]. When the Lindbladian dissipator is so structured, then the Hamiltonian part of the master equation becomes the canonical Hamiltonian. This canonical Hamiltonian contains information beyond just the system Hamiltonian about the part of the open system that evolves unitarily. A convenient way to find out the canonical Hamiltonian is given by
@@ III The central spin model and its exact dynamics
where $\omega_{0}$ is the transition frequency of the central spin-$1/2$ particle, $\epsilon$ is the interaction strength between the system and the bath with $\sqrt{N}$ being the scaling factor. $\omega/N$ is the scaled frequency of the bath. The evolution of the composite system SB for an arbitrary initial state $\rho_{SB}(0)=\rho_{S}(0)\otimes\rho_{B}(0)$, assuming separable initial condition, of the total system (where $\rho_{S}(0)$ and $\rho_{B}(0)$ are the initial states of the central spin and the bath, respectively) by means of the global unitary $U=e^{-iHt}$ is given by
@@
where $H$ is the total Hamiltonian. Here, we consider the Gibbs state as the initial state of the bath, which is given (using the spectral decomposition of the bath Hamiltonian $H_{B}=\frac{\omega}{N}J_{z}=\frac{\omega}{N}\sum_{n=0}^{N}\left\{\left(\frac{N}{2}-n\right)\ket{n}\bra{n}\right\}=\frac{\omega}{2}\sum_{n=0}^{N}\left\{\left(1-\frac{2n}{N}\right)\ket{n}\bra{n}\right\}$ | }) by
$\rho_{B}(0)$ $=\frac{e^{-\beta H_{B}}}{Z}=\frac{1}{Z}\sum_{n=0}^{N}e^{-\frac{\beta\omega}{2}\left(1-\frac{2n}{N}\right)}\ket{n}\bra{n},$ (32)
where $\beta=1/T$, |$\ket{n}$ is the standard computational basis, and $Z=\sum_{n=0}^{N}e^{-\frac{\beta\omega}{2}\left(1-\frac{2n}{N}\right)}$. To find out the exact dynamics of the system, we derive the spectral decomposition of the total Hamiltonian $H$. The nature of the eigenvalues and eigenvectors of the total Hamiltonian of the central spin model is illustrated in Appendix A. We use these eigenvalues and eigenvectors in Eq. (31) to obtain the exact dynamics of the central spin system. Considering an arbitrary initial state of the system $\rho_{S}(0)=\begin{pmatrix}\rho_{00}(0)&&\rho_{01}(0)\\ \rho_{10}(0)&&\rho_{11}(0)\end{pmatrix}$, the density matrix dictating the evolution of the central spin is given by
@@
$\delta^{*}(t)$ $=\frac{1}{Z}\left\{\frac{1}{1+\chi_{+}(1)^{2}}\left[e^{-i\left\{\lambda_{+}(1)-\frac{(\omega+\omega_{0})}{2}\right\}t-\frac{\beta\omega}{2}}+\chi_{+}(N)^{2}e^{-i\left\{\lambda_{+}(N)-\frac{(\omega+\omega_{0})}{2}\right\}t+\frac{\beta\omega}{2}}\right]+\frac{1}{1+\chi_{-}(1)^{2}}\left[e^{-i\left\{\lambda_{-}(1)-\frac{(\omega+\omega_{0})}{2}\right\}t-\frac{\beta\omega}{2}}+\chi_{-}(N)^{2}e^{-i\left\{\lambda_{-}(N)-\frac{(\omega+\omega_{0})}{2}\right\}t+\frac{\beta\omega}{2}}\right]\right\}$
$+\frac{1}{Z}\sum^{N}_{n=2}e^{-\frac{\beta\omega}{2}\left[1-\frac{2(n-1)}{N}\right]}\left\{\frac{e^{-i\lambda_{+}(n)t}}{1+\chi_{+}(n)^{2}}+\frac{e^{-i\lambda_{-}(n)t}}{1+\chi_{-}(n)^{2}}+e^{-i\lambda_{+}(n-1)t}\frac{\chi_{+}(n-1)^{2}}{1+\chi_{+}(n-1)^{2}}+e^{-i\lambda_{-}(n-1)t}\frac{\chi_{-}(n-1)^{2}}{1+\chi_{-}(n-1)^{2}}\right\}.$ (36)
Here, $\chi_{\pm}(n),b$ and $a_{n}$ comes from the structure of the eigenvectors of the total Hamiltonian $H$ discussed in Appendix A. The dynamical map $\Phi(t)$ of the system in the superoperator space is given by
@@
Figure 1: Variation of the central spin system’s density matrix element $\rho_{00}(t)$ with time for different values of the number of spins $N$ in the spin bath. Subplot (a) has $\epsilon=0.1,T=1$, (b) has $\epsilon=0.5,T=1$, (c) has $\epsilon=0.2,T=10$, and (d) has $\epsilon=0.2,T=0.1$. Further, $\omega_{0}=2.5$ and $\omega=2.0$ for all the plots. The system is initially taken to be in the excited state (|$\ket{\psi(0)}_{S}=\ket{0}_{S}$).
@@
Having discussed the dynamics of the central spin system, we now move on to study its thermodynamical aspects in the subsequent section.
Figure 2: Variation of (a) change in the internal energy of the system $\Delta U_{S}$, (b) change in the energy of the bath $Q_{B}$, (c) Work done by the bath $W$, and (d) entropy production $\Sigma$ with time $t$. The parameters are chosen to be $N=50$, $\omega=3$, $\omega_{0}=3.25$, and $T=0.25$. Further, the system’s initial state is taken to be $\frac{1}{2}\ket{0}+\frac{\sqrt{3}}{2}\ket{1}$.
@@ IV.1 Entropy production, heat, work and the laws of quantum thermodynamics
This state is used to calculate the change in energy of the bath $Q_{B}$, and then work done $W$ by the bath is calculated using $W=\Delta U_{S}+Q_{B}$, where $\Delta U_{S}$ is obtained from Eq. (41).
The variations of the above-mentioned thermodynamic quantities with time are plotted in Fig. 2. We observe that the change in the internal energy of the system $\Delta U_{S}$ is positive and undergoes revivals in time, as depicted by the oscillations. For higher values of system-bath interaction parameter $\epsilon$, we observe well-defined wavepacket-like structures. Further, at a very low value of $\epsilon$, the system’s internal energy change is approximately zero. The above observation matches the variations in the change in the energy of the bath $Q_{B}$ and the work done by the bath $W$. The values of the $Q_{B}$ are approximately zero for weak coupling ($\epsilon/\omega_{0}\approx 0.03$), which is consistent with the GKLS condition, and are otherwise negative, denoting that the energy of the bath decreases during the dynamics, and the positive values of $W$ denote that the bath works on the system. This observation is consistent with the first law of thermodynamics; that is, the energy of the bath decreases while it works on the system. In the process, the internal energy of the system increases. Interestingly, from Fig. 2(c), we observe that the amount of work required to turn off the interaction is approximately zero in the case of weak coupling ($\epsilon/\omega_{0}\approx 0.03$) and increases with an increase in the coupling strength. This observation is consistent with the literature [65].
@@ IV.2 HMF and thermodynamic entropy of the central spin model
In Fig. 4(a), we show the impact of interaction strength and bath temperature on the HMF (note that here we have taken a large spin bath $N=80$) by plotting the Hilbert-Schmidt norm $\left(\|O\|_{HS}=\sqrt{{\rm Tr}[O^{\dagger}O]}\right)$ of the difference between the system’s bare Hamiltonian $H_{S}$ and HMF $\mathcal{H}_{S}^{*}(\beta)$ as a function of inverse temperature $\beta$ and for different system-bath interaction strengths $\epsilon$. It can be observed from the plots that the HMF becomes equal to the system’s bare Hamiltonian at high temperatures and in the weak coupling regime. The difference is higher for higher $\epsilon$ and lower temperatures. However, the difference saturates as we keep on decreasing the temperature. The thermodynamic entropy of the system is plotted in Fig. 4(b) for different $\epsilon$ as a function of inverse temperature $\beta$. The thermodynamic entropy takes higher values at low temperatures when the system-bath coupling is higher. As we decrease the system-bath coupling, the entropy modulates more slowly for low-temperature values. At very low temperatures, the entropy can be seen to become zero, consistent with the third law of thermodynamics.
@@ IV.3 The canonical Hamiltonian of the central spin model
$\Im(\cdot)$ denotes the imaginary part of ()($\cdot$)( ⋅ ). Further, $\Omega(0)=\frac{\omega_{0}}{2}$.
@@ IV.3.1 Role of the canonical Hamiltonian in central spin’s thermodynamics
[…]
@@
The thermodynamic quantities computed in this section using the canonical Hamiltonian are plotted in Fig. 6, along with the corresponding quantities from the previous section to benchmark their variations. From an immediate inspection, we see that the quantities calculated using the canonical Hamiltonian show enigmatic behavior. The entropy production, Fig. 6(c), can be observed to be negative when computed using Eq. (54), violating the second law of thermodynamics. This equation of entropy production has two problems. First, it is calculated in the system’s subspace, albeit considering the canonical Hamiltonian carries the effect of the bath even after it is traced out. A similar observation of this expression of entropy production being negative was also observed in [91]. The other issue is that even if we replace the canonical Hamiltonian with the system’s Hamiltonian and further take it to be time-independent, the expression for the entropy production $\Sigma$ from Eq. (6) doesn’t match Eq. (54) due to the presence of work done by the bath $W$. Furthermore, it is observed from Fig. 6(c) that the entropy production $\Sigma$ in the form of Eq. (6) preserves the second law of thermodynamics ($\Sigma\geq 0$) for the same set of parameters. We also plot the quantity $\Sigma^{\prime}$, Eq. (61), where we replace the internal energy change of the bath $Q_{B}$ in Eq. (6) by the internal energy change of the bath due to time-dependent bath Hamiltonian $\Delta\widetilde{U}_{B}(t)$. However, this quantity also goes negative at various points in time. We refrain from calling it an expression of entropy production because even if we consider a time-dependent bath Hamiltonian, the entropy production [Eq. (5)], for the initial thermal state of the bath, results in Eq. (6) and not in Eq. (61).
@@
The above ambiguities resolve when we consider the original time-independent system Hamiltonian $H_{S}$, Eq. (III), for calculating the expectation values of the thermodynamic quantities. In this case, as observed in the previous section, the system gains energy while the bath loses energy, and the mismatch between them is the work done by the bath on the system, which is positive. This scenario is consistent with the first law of thermodynamics. Further, on considering $H_{S}$ as the system Hamiltonian during the dynamics, we get $\delta\widetilde{W}_{S}=0$, and $\delta\widetilde{Q}_{S}(t)=\Delta\widetilde{U}_{S}(t)=\Delta U_{S}$. Therefore, the ambiguity regarding the work done by the bath also resolves and is given by Eq. (14). Moreover, using the time-independent system and bath Hamiltonian $H_{S}$ and $H_{B}$, respectively, the entropy production is also positive, in consonance with Eq. (6) that considers irreversibility arising due to tracing out the bath. This scenario is consistent with the second law of thermodynamics.

### v3

@@ I Introduction
In this work, we study the strong-coupling non-Markovian quantum thermodynamics of a central spin interacting with a spin bath [73, 74, 75, 76, 77]. The baths can be broadly placed into the following categories [78]: (a) Bosonic, (b) fermionic, and (c) spin baths. The spin-boson model [8] and the Caldeira-Leggett model [11] are well-known examples of bosonic baths. The dynamics of these types of models have been studied in the literature [1, 2, 3]. Conducting electrons in the Anderson and the Kondo models are examples of a fermionic bath system. Numerically exact methods, for example, the fermionic hierarchical equation of motion, have been used to obtain the dynamics of the above systems [79, 80]. Further, the exact dynamics of noninteracting fermions (bosons) linearly coupled to thermal environments of noninteracting fermions can be found in [81]. Conversely, when dealing with spin baths [82], frequent use is made of perturbative methods or time-dependent quantum master equations [83, 84, 85]. The dissipative equation of motion approach has also been used to find the exact dynamics of the systems surrounded by a spin bath [86]. Studying these systems is crucial for understanding magnetic systems [87], quantum spin glasses [88], quantum batteries [89], NV-center [90], and superconducting systems [83], among others. Moreover, the finite nature of the bath renders the exact dynamics of the central spin model analytically tractable across arbitrary system-bath coupling regimes. This unique feature provides profound advantages for probing the non-Markovian effects on quantum thermodynamics, particularly in strong coupling regimes.
@@
Here, we develop the exact dynamics of the central spin model. This is then used to understand the quantum thermodynamic properties of the central spin. To this effect, we calculate the thermodynamic quantities of this system using a number of methods outlined above. We begin by analyzing entropy production, heat, and work while discussing the first and second laws of thermodynamics. We also find out the cost of turning the system and bath interactions on or off, and we explore a variation of the second law in finite bath models. Next, we calculate the thermal equilibrium state and entropy under strong coupling using the HMF technique, which aids in examining the third law of quantum thermodynamics. We then discuss the use of the canonical Hamiltonian as a replacement for the system’s bare Hamiltonian to calculate the thermodynamic quantities. We aim to arrive at the correct forms of the quantities, such as heat, work, internal energy, entropy production, and thermal equilibrium state in strongly coupled non-Markovian quantum systems and the correct forms of the corresponding laws of quantum thermodynamics, taking the central spin model as an example. Finally, we demonstrate the central spin model as a quantum battery, highlighting how a finite spin bath uniquely charges the central spin quantum battery. To summarize, we first present a brief discussion of the tools for handling strong-coupling non-Markovian quantum thermodynamics (Sec. II). This is followed in the later sections with new results on the exact dynamics of the central spin model and its quantum thermodynamics in the strong coupling non-Markovian regime, by applying the above tools.
@@ II Tools for exploring non-Markovian strong-coupling quantum thermodynamics
Here, we briefly dive into the key tools for probing quantum thermodynamics in a strongly coupled non-Markovian quantum system. We initiate the discussion from entropy production for a general quantum system surrounded by a bath (using this, we discuss the second law of quantum thermodynamics), followed by the discussion of the change in the system’s and bath’s internal energy and the work done as the mismatch between them. This leads us to the first law of quantum thermodynamics in a general system-bath scenario. In the next step, we discuss how the work done is related to the cost of turning the system-bath interactions on or off. To analyze the thermal equilibrium between the system and the bath in the strong-coupling regime, we discuss the formalism of the Hamiltonian of mean force and the mean force Gibbs state. Further, we discuss the canonical Hamiltonian of a system.
@@ II.1 Entropy production, heat and work
Considering the global system-bath dynamics for arbitrary initial states of the system and the bath [Eq. (2)], a formulation for the entropy production was developed in [91, 28]. Here, the dynamical evolution of the system is obtained by tracing out the bath, which is the origin of the irreversibility. The entropy production accounts for the contribution of two processes, one where we discard any information stored locally in the state of the bath, and the other where we discard the non-local information shared between the system and the bath, and is given by
@@
$Q_{B}$ $=\int_{0}^{t}d\tau\left\{-\dot{Q}_{S}(\tau)-\frac{d}{d\tau}\braket{V(\tau)}\right\}$
$=-\int_{0}^{t}d\tau\left\{i\left\langle[V(\tau),H_{S}(\tau)]\right\rangle\right\}+\braket{V(0)}-\braket{V(t)}.$ (13)
@@ II.3 The theory of canonical Hamiltonian
on the Lindblad (jump) and Hamiltonian operators, respectively. It leaves the structure $i[H_{S}(t),\rho_{S}]+\mathcal{D}_{t}(\rho_{S})$ intact. To get rid of this ambiguity and to render a unique dissipator, the jump operators $L_{j}^{\prime}$s need to be traceless, as shown in [98]. When the Lindbladian dissipator is so structured, then the Hamiltonian part of the master equation becomes the canonical Hamiltonian. This canonical Hamiltonian contains information beyond just the system Hamiltonian about the part of the open system that evolves unitarily. A convenient way to find out the canonical Hamiltonian is given by
@@ III The central spin model and its exact dynamics
where $\omega_{0}$ is the transition frequency of the central spin-$1/2$ particle, $\epsilon$ is the interaction strength between the system and the bath, with $\sqrt{N}$ being the scaling factor. $\omega/N$ is the scaled frequency of the bath. The evolution of the composite system SB for an arbitrary initial state $\rho_{SB}(0)=\rho_{S}(0)\otimes\rho_{B}(0)$, assuming separable initial condition, of the total system (where $\rho_{S}(0)$ and $\rho_{B}(0)$ are the initial states of the central spin and the bath, respectively) by means of the global unitary $U=e^{-iHt}$ is given by
@@
[…]
$\rho_{B}(0)$ $|,\displaystyle=\frac{e^{-\beta H_{B}}}{Z}=\frac{1}{Z}\sum_{n=0}^{N}e^{-\frac{\beta\omega}{2}\left(1-\frac{2n}{N}\right)}\ket{n}\bra{n}$, (32)
[…]
@@
$\delta^{*}(t)$ $=\frac{1}{Z}\left\{\frac{1}{1+\chi_{+}(1)^{2}}\left[e^{-i\left\{\lambda_{+}(1)-\frac{(\omega+\omega_{0})}{2}\right\}t-\frac{\beta\omega}{2}}+\chi_{+}(N)^{2}e^{i\left\{\lambda_{+}(N)-\frac{(\omega+\omega_{0})}{2}\right\}t+\frac{\beta\omega}{2}}\right]+\frac{1}{1+\chi_{-}(1)^{2}}\left[e^{-i\left\{\lambda_{-}(1)-\frac{(\omega+\omega_{0})}{2}\right\}t-\frac{\beta\omega}{2}}+\chi_{-}(N)^{2}e^{i\left\{\lambda_{-}(N)-\frac{(\omega+\omega_{0})}{2}\right\}t+\frac{\beta\omega}{2}}\right]\right\}$
$+\frac{1}{Z}\sum^{N}_{n=2}e^{-\frac{\beta\omega}{2}\left[1-\frac{2(n-1)}{N}\right]}\left\{\left(\frac{e^{-i\lambda_{+}(n)t}}{1+\chi_{+}(n)^{2}}+\frac{e^{-i\lambda_{-}(n)t}}{1+\chi_{-}(n)^{2}}\right)\times\left(e^{i\lambda_{+}(n-1)t}\frac{\chi_{+}(n-1)^{2}}{1+\chi_{+}(n-1)^{2}}+e^{i\lambda_{-}(n-1)t}\frac{\chi_{-}(n-1)^{2}}{1+\chi_{-}(n-1)^{2}}\right)\right\}.$ (36)
Here, $\chi_{\pm}(n),b$ and $a_{n}$ come from the structure of the eigenvectors of the total Hamiltonian $H$ discussed in Appendix A. The dynamical map $\Phi(t)$ of the system in the superoperator space is given by
@@
[…]
@@
Having discussed the dynamics of the central spin system, we now move on to study its thermodynamic aspects in the subsequent section.
[…]
@@ IV.1 Entropy production, heat, work and the laws of quantum thermodynamics
This state is used to calculate the change in energy of the bath $Q_{B}$, and then the work done $W$ by the bath is calculated using $W=\Delta U_{S}+Q_{B}$, where $\Delta U_{S}$ is obtained from Eq. (41).
The variations of the above-mentioned thermodynamic quantities with time are plotted in Fig. 2. We observe that the change in the internal energy of the system $\Delta U_{S}$ is positive and undergoes revivals in time, as depicted by the oscillations. For higher values of the system-bath interaction parameter $\epsilon$, we observe well-defined wavepacket-like structures. Further, at a very low value of $\epsilon$, the system’s internal energy change is approximately zero. The above observation matches the variations in the change in the energy of the bath $Q_{B}$ and the work done by the bath $W$. The values of the $Q_{B}$ are approximately zero for weak coupling ($\epsilon/\omega_{0}\approx 0.03$), which is consistent with the GKLS condition, and are otherwise negative, denoting that the energy of the bath decreases during the dynamics, and the positive values of $W$ denote that the bath works on the system. This observation is consistent with the first law of thermodynamics; that is, the energy of the bath decreases while it works on the system. In the process, the internal energy of the system increases. Interestingly, from Fig. 2(c), we observe that the amount of work required to turn off the interaction is approximately zero in the case of weak coupling ($\epsilon/\omega_{0}\approx 0.03$) and increases with an increase in the coupling strength. This observation is consistent with the literature [65].
@@ IV.2 HMF and thermodynamic entropy of the central spin model
In Fig. 4(a), we show the impact of interaction strength and bath temperature on the HMF (note that here we have taken a large spin bath $N=80$) by plotting the Hilbert-Schmidt norm $\left(\|O\|_{HS}=\sqrt{{\rm Tr}[O^{\dagger}O]}\right)$ of the difference between the system’s bare Hamiltonian $H_{S}$ and HMF $\mathcal{H}_{S}^{*}(\beta)$ as a function of inverse temperature $\beta$ and for different system-bath interaction strengths $\epsilon$. It can be observed from the plots that the HMF becomes equal to the system’s bare Hamiltonian at high temperatures and in the weak coupling regime. The difference is higher for higher $\epsilon$ and lower temperatures. However, the difference saturates as we keep decreasing the temperature. The thermodynamic entropy of the system is plotted in Fig. 4(b) for different $\epsilon$ as a function of inverse temperature $\beta$. The thermodynamic entropy takes higher values at low temperatures when the system-bath coupling is higher. As we decrease the system-bath coupling, the entropy modulates more slowly for low-temperature values. At very low temperatures, the entropy can be seen to become zero, consistent with the third law of thermodynamics.
@@ IV.3 The canonical Hamiltonian of the central spin model
$\Im(\cdot)$ denotes the imaginary part of $(\cdot)$. Further, $\Omega(0)=\frac{\omega_{0}}{2}$.
@@ IV.3.1 Role of the canonical Hamiltonian in central spin’s thermodynamics
$\delta\widetilde{W}_{S}(t)$ $=\int_{0}^{t}d\tau~~{\rm Tr}\left[\dot{H}^{\rm can}_{S}(\tau)\rho_{S}(\tau)\right],$
$\delta\widetilde{Q}_{S}(t)$ $=\int_{0}^{t}d\tau~~{\rm Tr}\left[H^{\rm can}_{S}(\tau)\dot{\rho}_{S}(\tau)\right].$ (53)
@@
The thermodynamic quantities computed in this section using the canonical Hamiltonian are plotted in Fig. 6, along with the corresponding quantities from the previous section to benchmark their variations. From an immediate inspection, we see that the quantities calculated using the canonical Hamiltonian show enigmatic behavior. The entropy production, Fig. 6(c), can be observed to be negative when computed using Eq. (54), violating the second law of thermodynamics. This equation of entropy production has two problems. First, it is calculated in the system’s subspace, albeit considering the canonical Hamiltonian carries the effect of the bath even after it is traced out. A similar observation of this expression of entropy production being negative was also observed in [91]. The other issue is that even if we replace the canonical Hamiltonian with the system’s Hamiltonian and further take it to be time-independent, the expression for the entropy production $\Sigma$ from Eq. (6) doesn’t match Eq. (54) due to the presence of work done by the bath $W$. Furthermore, it is observed from Fig. 6(c) that the entropy production $\Sigma$ in the form of Eq. (6) preserves the second law of thermodynamics ($\Sigma\geq 0$) for the same set of parameters. We also plot the quantity $\Sigma^{\prime}$, Eq. (61), where we replace the internal energy change of the bath $Q_{B}$ in Eq. (6) by the internal energy change of the bath due to the time-dependent bath Hamiltonian $\Delta\widetilde{U}_{B}(t)$. However, this quantity also goes negative at various points in time. We refrain from calling it an expression of entropy production because even if we consider a time-dependent bath Hamiltonian, the entropy production [Eq. (5)], for the initial thermal state of the bath, results in Eq. (6) and not in Eq. (61).
@@
The above ambiguities resolve when we consider the original time-independent system Hamiltonian $H_{S}$, Eq. (III), for calculating the expectation values of the thermodynamic quantities. In this case, as observed in the previous section, the system gains energy while the bath loses energy, and the mismatch between them is the work done by the bath on the system, which is positive. This scenario is consistent with the first law of thermodynamics. Further, on considering $H_{S}$ as the system Hamiltonian during the dynamics, we get $\delta\widetilde{W}_{S}=0$, and $\delta\widetilde{Q}_{S}(t)=\Delta\widetilde{U}_{S}(t)=\Delta U_{S}$. Therefore, the ambiguity regarding the work done by the bath also resolves and is given by Eq. (14). Moreover, using the time-independent system and bath Hamiltonians $H_{S}$ and $H_{B}$, respectively, the entropy production is also positive, in consonance with Eq. (6) that considers irreversibility arising due to tracing out the bath. This scenario is consistent with the second law of thermodynamics.